OpenStudy (anonymous):

According to the 2000 U.S. Census Bureau, 49.65% of Texas residents were male. Consider the experiment of selecting ten Texans at random. If you are interested in determining the probabilities of selecting x males, is this a binomial probability distribution? And if so, what are the values of n, p, and q?

OpenStudy (anonymous):

in a binomial probability, you only have two choices... in this case if it is not male then it is a female.... in the experiment, 10 Texans will be selected at random... therefore n = 10 trials since there are only two sex upon random selection... male or female, the probability P(male) = p = 1/2 = 0.5 P(female) = q = 1/2 = 0.5

Mercury:

mostly agree with the previous response, but want to point out that the probability is not 1/2 in this case since it's based on a Texas census, which shows a slightly different probability (49.65% of Texas residents were male.)